What is the Na+1 ion concentration (in moles per liter) when 0.0234 g Na2CO3 is dissolved in enough water to make 150. mL of solution?
1
Expert's answer
2012-12-24T11:22:51-0500
The molar mass of Na2CO3 is 106 g/mol. The amount of Na2CO3 is n(Na2CO3) = m(Na2CO3) / M(Na2CO3) = 0.0234 g / 106 g/mol = 0.00022 mol.
The molar concentration of Na2CO3 is c(Na2CO3) = n(Na2CO3) / V(solution) = 0.00022 mol / 0.150 L = 0.0015 M (moles/L)
According to dissociation reaction below, the concentration of Na(+1) ion is twice more of concentration of Na2CO3. c(Na+) = 2 * c(Na2CO3) = 0.003 M.
Na2CO3 = 2 Na(+) + CO3(2-)
Answer: 0.003 M